### v2

@@ 6.1 Analysis of retrieval
Based on the above analysis, we summarize two empirical rules as follows: 1) For a dataset with low or high intra-class visual variance, it is imperative to employ a retrieval model with correspondingly strong or weak capabilities to guarantee stable performance improvements across scenarios. 2) In low-shot settings, the inference outcomes are dominantly influenced by the retrieval results.
6.2 Tuning retrieval parameters
The final prediction distribution in RePrompt is influenced by the hyperparameters $\gamma$ and $\lambda$, and the prompt depth $J$. Ablation experiments are conducted to investigate the impact of these hyperparameters on the final outcomes across different datasets.
As illustrated in Figure 9, from top to bottom, ablation experiments are carried out on FGVC Aircraft, ImageNet, and Oxford Pets under the 16-shot learning setting. The experiments adhere to the default configurations, except for modifying the depth $J$, $\gamma$, and $\lambda$ parameters. The experiment data are averaged over the results of experiments on seeds 1,2,3. The reference configuration is marked by a gray vertical dotted line.
[…]
$\gamma$ varies. The hyperparameters $\gamma$ in the guiding loss $\mathcal{L}$ controls the intensity of the retrieval enhancement. As evidenced by the right part of Figure 9, the result continues to improve as $\gamma$ decreases. These results demonstrate that lower values of $\gamma$ enhance the reliance on retrieval-enhanced features, leading to significant improvements in classification accuracy. However, for FGVC Aircraft, the highest accuracy is observed at $\gamma=10^{-1}$, which is caused by its smallest intra-class visual variance.
In summary, the hyper-parameters associated with retrieval critically influence the efficacy of retrieval enhancement. These findings highlight the necessity of striking an optimal balance between incorporating valuable retrieved information and mitigating the potential introduction of noise during the learning process. We provide valuable insights that are instrumental for the design of related methods and establish a guiding framework for future research in this domain.
7 Conclusion
@@
In this paper, we investigate retrieval-enhanced visual prompt learning for visual-language models. Specifically, we propose integrating the results from a retrieval cache model into model inference using two approaches: prompts and predicted distributions. Additionally, We introduce a prompt learner that dynamically adapts prompts based on retrieval results and an auxiliary classification adapter. To regulate the degree of retrieval enhancement, we introduce the prior distribution obtained from semi-parametric retrieval into the cross-entropy loss to guide prompt tuning.
@@
8 Data Availability Statements
The datasets that support the findings of this study are openly available in
1) Few-shot image classification: ImageNet, Caltech101, OxfordPets, StanfordCars, Flowers102, Food101, FGVCAircraft, SUN397, DTD, EuroSAT, UCF101. Download split files from this link
2) Domain generalization task: ImageNetV2, ImageNet-Sketch, ImageNet-A, ImageNet-R.
3) Few-shot video understanding: HMDB51, UCF101, Something Something v2 (SSv2), Kinetics
4) Few-shot Multi-few image classification: MVImgNet
Acknowledgments
This work was supported by National Key R&D Program of China (No. 2022ZD0118700) and Natural Science Foundation of Zhejiang Province (NO. LY21F030018).

### v3

@@ VI-A Analysis of Retrieval
Based on the above analysis, we summarize two empirical rules as follows: 1) For a dataset with low or high intra-class visual variance, it is imperative to employ a retrieval model with correspondingly strong or weak capabilities to guarantee stable performance improvements across scenarios. For example, if the dataset has a large intra-class visual variation, we need to control the impact of the research branch on the results by limiting the number of retrieved samples and retrieval-enhanced prompt. 2) In low-shot settings, the inference outcomes are dominantly influenced by the retrieval results.
VII Conclusion
@@
In this paper, we propose RePrompt, a retrieval-based framework that enhances the performance of visual-language modes on few-shot classification tasks. Specifically, we integrate the retrieved results from a retrieval cache model into model inference using two approaches: prompts and predicted distributions. Additionally, We introduce a prompt learner that dynamically adapts prompts based on retrieval results and an auxiliary classification adapter. To regulate the degree of retrieval enhancement, we introduce the prior distribution obtained from semi-parametric retrieval into the cross-entropy loss to guide prompt tuning.